OpenStudy (anonymous):

About 55% of emergency room visits are unnecessary. Assuming each emergency room visit is independent. (Hints: first define the random variable X. Then write the objective of each question in terms of X.) a) Suppose the distribution of unnecessary emergency room visits can be approximated by a normal distribution. What is the probability that AT LEAST two of them need to be in the emergency room? sample size 15 patients @kropot72 ...

OpenStudy (anonymous):

dont know if u can help me @kropot72

OpenStudy (kropot72):

The binomial distribution applies to this question. Let the number of necessary visits be X. \[P(X=0)=\left(\begin{matrix}15 \\ 0\end{matrix}\right)0.45^{0}0.55^{15}=0.000127\] \[P(X=1)=\left(\begin{matrix}15 \\ 1\end{matrix}\right)0.45^{1}0.55^{14}=0.001565\] Therefore we can find the required probability as follows: \[P(X \ \ge2)=1.000000-[P(X=0)+P(X=1)]\]

OpenStudy (anonymous):

exactly wat i thought..thanks

OpenStudy (kropot72):

You're welcome :)